# Optimal Strategy Revision in Population Games:

### v1

@@ III Optimal Strategy Revision Protocol Design
We begin by establishing the relationship between the MFG system and the optimal strategy revision process in PG, as illustrated in Fig. 1. Lemma 1 shows the connection between the state distribution evolution in finite-state MFG and the ED model, followed by Theorem 1, which provides a solution to our main problem.
Lemma 1
The ED (2) is equivalent to the FK equation (9) if the strategy revision protocol satisfies
$\rho_{ij}(p(t),x(t))=\begin{cases}\alpha_{ij}(t)&\text{if}~{}i\neq j\\ 0&\text{otherwise}\end{cases},~{}i,j\in\mathbb{S}.$ (11)
Proof:
By definition, we have $\sum_{j\in\mathbb{S}\setminus\{i\}}\alpha_{ij}(t)=-\alpha_{ii}(t)$, which allows us to rewrite (9) as:
$\dot{x}_{i}(t)$ $=\textstyle\sum_{j\in\mathbb{S}\setminus\{i\}}\alpha_{ji}(t)x_{j}(t)+\alpha_{ii}(t)x_{i}(t).$
This simplifies further to:
$\dot{x}_{i}(t)\!=\!\textstyle\sum_{j\in\mathbb{S}\setminus\{i\}}x_{j}(t)\alpha_{ji}(t)\!-\!x_{i}(t)\sum_{j\in\mathbb{S}\setminus\{i\}}\alpha_{ij}(t).$ (12)
Notably, the value of $\alpha_{ii}(t)$ does not influence (12) because the sums can be extended to include all $j\in\mathbb{S}$ without altering the results, as the contributions for $j=i$ are null. With $\rho_{ij}(p(t),x(t))=\alpha_{ij}(t)$ for all $i\neq j$, we complete the proof. [^3] Also, in the context of ED, when considering pairwise comparison revision protocols, it holds that $\rho_{ii}(p(t),x(t))=0$, for all $i\in\mathbb{S}$, showing that (2) can be written using the same sum indexing as in (12), i.e., over $j\in\mathbb{S}\setminus\{i\}$.   $\blacksquare$
@@
$\dot{v}_{i}(t,x(t))$ $=-\frac{1}{2}\textstyle\sum_{j\in\mathbb{S}}\dfrac{[v_{j}(t,x(t))\!-\!v_{i}(t,x(t))]_{+}^{2}}{q_{ij}(t)}$
$\qquad-F_{i}(x(t))$ (13a)
@@
$\dot{x}_{i}(t)$ $=\textstyle\sum_{j\in\mathbb{S}}x_{j}(t)\dfrac{[v_{i}(t,x(t))-v_{j}(t,x(t))]_{+}}{q_{ji}(t)}$
$\quad-x_{i}(t)\textstyle\sum_{j\in\mathbb{S}}\dfrac{[v_{j}(t,x(t))-v_{i}(t,x(t))]_{+}}{q_{ij}(t)},$ (14a)
@@
where $x({t_{0}})\in\Delta$.
Proof:
We begin by determining the optimal state transition matrix $\alpha(t)$ within the MFG framework, where the finite time-horizon cost function in (6) is defined with $c(s(t),x(t),\alpha_{s(t)}(t))=\sum_{j\in\mathbb{S}}\frac{q_{s(t),j}(t)}{2}\alpha_{s(t),j}^{2}(t)-F_{s(t)}(x(t))$ and $\psi(s(T),x(T))=-F_{s(T)}(x(T))$. The Hamiltonian is given by:
$H_{i}(w(t,x(t)),x(t))=\textstyle\min_{\alpha_{i}\in\mathbb{R}_{i,0}^{n}}\bigg{(}\sum_{j\in\mathbb{S}}\frac{q_{ij}(t)}{2}\alpha_{ij}^{2}\\ -F_{i}(x(t))+\textstyle\sum_{j\in\mathbb{S}}\alpha_{ij}(w_{j}(t,x(t))-w_{i}(t,x(t)))\bigg{)}.$
$\alpha_{ij}(t)=\frac{[w_{i}(t,x(t))-w_{j}(t,x(t))]_{+}}{q_{ij}(t)},~{}\forall i\neq j.$ (15)
Consequently, we derive the Hamiltonian function as follows:
$H_{i}(w(t,x(t)),x(t))\\ =-\frac{1}{2}\textstyle\sum_{j\in\mathbb{S}}\dfrac{[w_{i}(t,x(t))-w_{j}(t,x(t))]_{+}^{2}}{q_{ij}(t)}-F_{i}(x(t)),$
for all $i$ in $\mathbb{S}$. Since $w_{i}(t,x)$ is the value function of (6), we define $v_{i}(t,x)=-w_{i}(t,x)$, making $v_{i}(t,x)$ the optimal cost-to-go for (4). By combining (7) with Lemma 1, we use $\alpha_{ij}(t)$ and $v_{i}(t,x(t))$ to define $\rho_{ij}(p(t),x(t))$ and $p_{i}(t)$, respectively. This concludes the proof. $\blacksquare$
As stated in Theorem 1, the payoff dynamics model (3) is given by:
$G_{i}(t,p(t),x(t))=-\frac{1}{2}\textstyle\sum_{j\in\mathbb{S}}\dfrac{[p_{j}(t)-p_{i}(t)]_{+}^{2}}{q_{ij}(t)}-F_{i}(x(t))$
with the terminal condition $p_{i}(T)=F_{i}(x(T))$ for all $i$ in $\mathbb{S}$.
@@ III-A Direct Connection with Conventional ED Models
The following corollary demonstrates how specific classes of existing ED models can be derived from Theorem 1 by identifying a corresponding weight $q_{ij}(t)$ in (4) for each class. Each resulting ED model is named after its counterpart in the literature, as detailed in Remark 1, where we provide an interpretation of the associated revision protocol. This interpretation offers valuable insights into the behaviors observed in PG.
@@
Consider the following choices for the weight $q_{ij}(t)$ in (4): $i)$ $q_{ij}(t)=1$, $ii)$ $q_{ij}(t)=1/x_{j}(t)$, and $iii)$ $q_{ij}(t)=x_{i}(t)$. For each case, the corresponding ED models are as follows:
$i)$ Smith dynamics and payoff dynamics:
@@
$ii)$ Replicator dynamics and payoff dynamics:
@@
$iii)$ Projection dynamics and payoff dynamics:
@@
All models satisfy the terminal condition $v_{i}(T,x(T))=F_{i}(x(T))$ and the initial condition $x(t_{0})\in\Delta$.
Corollary 1 highlights interesting connections between the choice of $q_{ij}(t)$ and several well-known properties in ED, as elaborated in the following remarks.
@@
$i)$ For the replicator dynamics, (4) diverges as the proportion of decision-makers approaches zero, i.e., $q_{ij}(t)\to\infty$ when $x_{j}(t)\to 0$. Consequently, choosing $q_{ij}(t)={1}/{x_{j}(t)}$ acts as a penalty function, ensuring that the boundary $\partial\Delta$ of the simplex is never reached. This aligns with the observation that under the replicator dynamics, the boundary $\partial\Delta$ acts as a repeller. This interpretation is consistent with the permanence theorem discussed in [11].
$ii)$ As $T\to t_{0}$ in (13), we recover $v_{i}(t_{0},x(t_{0}))=F_{i}(x(t_{0}))$. Substituting this result into (14) reproduces the standard ED models, as in (2), widely studied in the literature [6, 12, 13].
In the following corollary, we consider migration constraints that restrict agents from switching between specific pairs of strategies. These permissible strategy switches are represented by an undirected connected graph $(\mathbb{S},\mathbb{L},A)$, where $\mathbb{L}\subseteq\{(i,j)\,|\,i,j\in\mathbb{S}\}$ denotes the set of allowed migrations (graph links), and a symmetric matrix $A\in\{0,1\}^{n\times n}$ is the adjacency matrix. The entries $a_{ij}$ of $A$ capture the migration constraints, such that $(i,j)\in\mathbb{L}$ if and only if $a_{ij}=1$. The set of neighbors of strategy $i$ is defined as $\mathbb{S}_{i}=\{j\in\mathbb{S}\,|\,(j,i)\in\mathbb{L}\}$. Prior works [12, 13, 14] have demonstrated that migration constraints lead to distributed population dynamics by incorporating the adjacency matrix $A$ – or equivalently, the migration constraints – into the revision protocol.
@@
Consider the following choices for the weights $q_{ij}(t)$ in (4) using $A=(a_{ij})_{i,j\in\mathbb{S}}$: $i)$ $q_{ij}(t)={1}/{a_{ij}}$, $ii)$ $q_{ij}(t)={1}/{(a_{ij}x_{j}(t))}$, and $iii)$ $q_{ij}(t)={x_{i}(t)}/{a_{ij}}$. For each case, the corresponding ED models are as follows:
$i)$ Distributed Smith dynamics and payoff dynamics:
@@
$ii)$ Distributed replicator dynamics and payoff dynamics:
@@
$iii)$ Distributed projection dynamics and payoff dynamics:
@@
All models satisfy the terminal condition $v_{i}(T,x(T))=F_{i}(x(T))$ and the initial condition $x(t_{0})\in\Delta$.
@@ III-B Positive Correlation, Nash Stationarity, and Convergence
The payoff dynamics model (13) and the ED model (14) satisfy the following positive correlation: $V(p(t),x(t))\neq\mathbf{0}$ implies $p^{\top}(t)V(p(t),x(t))>0$, where $p(t)=v(t,x(t))$, for all $t$ in $[t_{0},T]$.
@@
$=\textstyle\sum_{i\in\mathbb{S}}\sum_{j\in\mathbb{S}}x_{j}(t)\dfrac{[v_{i}(t,x(t))-v_{j}(t,x(t))]_{+}^{2}}{q_{ji}(t)}.$
@@
Consider the modulo-addition stationary solution of the associated MFG system, as explained in (10). The stationary solution for (13) and (14) is given by: $v(t,\bar{x})=\kappa(t-t_{0})\mathbf{1}_{n}+v(t_{0},\bar{x}),~{}\forall t\in[t_{0},T]$, where $\bar{x}$ is a Nash equilibrium of $F$, and $v(T,\bar{x})$ is its associated payoff vector given by $v(T,\bar{x})=F(\bar{x})$.
@@
$\dot{v}(t,\bar{x})=\kappa\mathbf{1}_{n}\implies v(t,\bar{x})=\kappa(t-t_{0})\mathbf{1}_{n}+v(t_{0},\bar{x}),$
with $v(T,\bar{x})=F(\bar{x})$. From (10b), it follows that $V(p(t),\bar{x})=\mathbf{0}$. Following the same reasoning as in the proof of Proposition 1, this condition implies $\bar{x}_{i}[F_{j}(\bar{x})-F_{i}(\bar{x})]_{+}=0,~{}\forall i,j\in\mathbb{S}$, thereby establishing that $\bar{x}$ is a Nash equilibrium.  $\blacksquare$
In the following corollary, we provide a convergence result for a certain class of contractive population games.
Corollary 3
Suppose a population game $F$ is contractive and satisfies
$\left(F(x)-F(y)\right)^{\top}(x-y)\leq-\epsilon\|x-y\|_{2}^{2},$ (19)
where $\epsilon$ is a positive constant. Then, $x(t)$ converges to a Nash equilibrium of $F$.
Proof:
By Proposition 2, the modulo-addition stationary solution corresponds to a Nash equilibrium $x^{\ast}$. Invoking [7, Theorem 3], which states that $x(t)$ converges to the modulo-addition stationary solution when $F$ satisfies (19), we conclude that $x(t)$ converges to $x^{\ast}$. $\blacksquare$
@@
For any contractive population game $F$, modifying $F$ by defining $F_{\epsilon}(x)=F(x)-\epsilon\,x$ with any positive $\epsilon$ ensures that $F_{\epsilon}$ satisfies (19). Let $x^{\ast}(\epsilon)$ denote the Nash equilibrium of $F_{\epsilon}$, which also serves as the $\epsilon$-Nash equilibrium of $F$. If $F$ is a continuous function of $x$, it follows that:
@@
where $x^{\ast}$ is the Nash equilibrium of $F$.
@@ IV Illustrative Examples
Data: $F$, $x_{0})\in\Delta$, $t_{0},T\in\mathbb{R}_{+}$ with $T>t_{0}$, $N\in\mathbb{Z}_{+}$, $\varepsilon_{f}\in\mathbb{R}_{+}$
@@
10       $x^{(k+1)}_{[t_{0},T]}\leftarrow(1-\alpha)x^{(k)}_{[t_{0},T]}+\alpha x^{\prime}_{[t_{0},T]}$
11       $e\leftarrow\int_{t_{0}}^{T}\|x^{(k)}(t^{\prime})-x^{(k+1)}(t^{\prime})\|_{2}^{2}\,\mathrm{d}t^{\prime}$triangleright To check if the fixed-point is reached
@@ 12       $k\leftarrow k+1$
We provide numerical examples to illustrate the main results of this paper. To compute the solutions for (13) and (14), we use Algorithm 1, adapted from the literature [15]. The algorithm is specifically designed to find the fixed point of (13) and (14).
@@
We consider the following two population games.
@@
$F(x)=\begin{pmatrix}F_{1}(x)\\ F_{2}(x)\\ F_{3}(x)\end{pmatrix}=-\begin{pmatrix}3x_{1}+x_{3}\\ 2x_{2}+x_{3}\\ x_{1}+x_{2}+3x_{3}\end{pmatrix}.$ (20)
For (4), we set $q_{ij}=1,\forall i,j\in\mathbb{S}$, $t_{0}=0$, and $T=6$. We apply Algorithm 1 with $\alpha=0.01$, $N=100$, and $\varepsilon_{f}=0$.
@@
For (4), we set $q_{ij}=1,\forall i,j\in\mathbb{S}$, $t_{0}=0$, and $T=6$. We apply Algorithm 1 with $\alpha=0.001$, $N=6000$, and $\varepsilon_{f}=0$.
@@
Fig. 3 compares the trajectories generated by the optimal strategy revision protocol with those obtained using the Smith ED model in the RPS game (21). As shown, the population state trajectory determined by the optimal protocol exhibits reduced oscillations and converges more quickly to the Nash equilibrium of (21).

### v2

@@ III Optimal Strategy Revision Protocol Design
We begin by establishing the relationship between the MFG system and the optimal strategy revision process in PG, as illustrated in Fig. 1. The following theorem establishes the connection between the state distribution dynamics in finite-state MFGs and the ED model, providing a solution to our main problem. The proof is presented in Appendix -A. We note that the existence and uniqueness of solutions to the main problem are ensured by the results established for (6), as discussed in §II-B.
@@
$\dot{v}_{i}(t,x(t))$ $=-\frac{1}{2}\sum_{j\in\mathbb{S}}\dfrac{[v_{j}(t,x(t))\!-\!v_{i}(t,x(t))]_{+}^{2}}{q_{ij}(x(t))}\!-\!F_{i}(x(t))$ (11a)
@@
$\dot{x}_{i}(t)$ $=\textstyle\sum_{j\in\mathbb{S}}x_{j}(t)\dfrac{[v_{i}(t,x(t))-v_{j}(t,x(t))]_{+}}{q_{ji}(x(t))}$
$\qquad-x_{i}(t)\textstyle\sum_{j\in\mathbb{S}}\dfrac{[v_{j}(t,x(t))-v_{i}(t,x(t))]_{+}}{q_{ij}(x(t))},$ (12a)
@@
where $x_{0}\in\Delta$. $\square$
As stated in the theorem, the payoff dynamics model (3) is given by:
$G_{i}(p(t),x(t))=-\frac{1}{2}\textstyle\sum_{j\in\mathbb{S}}\dfrac{[p_{j}(t)-p_{i}(t)]_{+}^{2}}{q_{ij}(x(t))}-F_{i}(x(t))$
with the terminal condition $p_{i}(T)=F_{i}(x(T))$ for all $i\in\mathbb{S}$.
@@ III-A Direct Connection with Conventional ED Models
The following corollary demonstrates how specific classes of existing ED models can be derived from Theorem 1 by identifying a corresponding weight $q_{ij}(x(t))$ in (4) for each class. Each resulting ED model is named after its counterpart in the literature, as discussed in Remark 1, where we provide an interpretation of the associated revision protocol. This interpretation offers valuable insights into the behaviors observed in PG. For brevity, we restrict attention to the case where $x(t)\in\mathrm{int}(\Delta)$ for all $t\in[t_{0},T]$.
@@
Consider the following choices for the weight $q_{ij}(x(t))$ in (4): $i)$ $q_{ij}(x(t))=1$, ii)ii) $q_{ij}(x(t))=1/x_{j}(t)$, and $iii)$ $q_{ij}(x(t))=x_{i}(t)$. For each case, the corresponding ED models are as follows:
$i)$ Smith dynamics and payoff dynamics ($q_{ij}(x(t))=1$):
@@
ii)ii) Replicator dynamics and payoff dynamics ($q_{ij}(x(t))=1/x_{j}(t)$):
@@
$iii)$ Projection dynamics and payoff dynamics ($q_{ij}(x(t))=x_{i}(t)$):
@@
All models satisfy the terminal condition $v_{i}(T,x(T))=F_{i}(x(T))$ and the initial condition $x(t_{0})=x_{0}\in\mathrm{int}(\Delta)$. $\square$
Corollary 1 highlights interesting connections between the choice of $q_{ij}(x(t))$ and several well-known properties in ED, as elaborated in the following remarks.
@@
$i)$ For the replicator dynamics, when $\rho_{ij}(p(t),x(t))>0$, (4) may diverge as the proportion of decision-makers approaches zero, i.e., $q_{ij}(x(t))\to\infty$ when $x_{j}(t)\to 0$. Consequently, choosing $q_{ij}(x(t))={1}/{x_{j}(t)}$ serves as a penalty function that ensures the boundary $\partial\Delta$ of the simplex remains invariant. This behavior is consistent with the well-known property that, under replicator dynamics, the boundary $\partial\Delta$ forms an invariant set [Hofbauer_1987].
ii)ii) As $T\to t_{0}$ in (11), we recover $v_{i}(t_{0},x(t_{0}))=F_{i}(x(t_{0}))$. Substituting this result into (12) reproduces the standard ED models, as in (2), widely studied in the literature [Sandholm2010-SANPGA-2, BaObQu2017, EDM_constraints]. In contrast, when $T>t_{0}$, the presence of payoff dynamics—rather than static payoff functions—allows agents to optimize based on their cost-to-go, thereby exhibiting forward-looking strategic decision-making. $\square$
In the following corollary, we consider migration constraints that restrict agents from switching between specific pairs of strategies. These permissible strategy switches are represented by an undirected connected graph $(\mathbb{S},\mathbb{L},A)$, where $\mathbb{L}\subseteq\{(i,j)\,|\,i,j\in\mathbb{S}\}$ denotes the set of allowed migrations (graph links), and a symmetric matrix $A\in\{0,1\}^{n\times n}$ is the adjacency matrix. The entries $a_{ij}$ of $A$ capture the migration constraints, such that $(i,j)\in\mathbb{L}$ if and only if $a_{ij}=1$. The set of neighbors of strategy $i$ is defined as $\mathbb{S}_{i}=\{j\in\mathbb{S}\,|\,(j,i)\in\mathbb{L}\}$. Prior works [BaObQu2017, EDM_constraints, Tembine_2008] have demonstrated that migration constraints lead to distributed population dynamics by incorporating the adjacency matrix $A$ – or equivalently, the migration constraints – into the revision protocol.
@@
Consider the following choices for the weights $q_{ij}(x(t))$ in (4) using $A=(a_{ij})_{i,j\in\mathbb{S}}$: $i)$ $q_{ij}(x(t))={1}/{a_{ij}}$, ii)ii) $q_{ij}(x(t))={1}/{(a_{ij}x_{j}(t))}$, and $iii)$ $q_{ij}(x(t))={x_{i}(t)}/{a_{ij}}$. For each case, the corresponding ED models are as follows:
$i)$ Distributed Smith dynamics and payoff dynamics ($q_{ij}(x(t))=1/a_{ij}$):
@@
ii)ii) Distributed replicator dynamics and payoff dynamics ($q_{ij}(x(t))=1/(a_{ij}x_{j}(t))$):
@@
$iii)$ Distributed projection dynamics and payoff dynamics ($q_{ij}(x(t))=x_{i}(t)/a_{ij}$):
@@
All models satisfy the terminal condition $v_{i}(T,x(T))=F_{i}(x(T))$ and the initial condition $x(t_{0})=x_{0}\in\mathrm{int}(\Delta)$. $\square$
@@ III-B Positive Correlation, Nash Stationarity, and Convergence
The payoff dynamics model (11) and the ED model (12) satisfy the following positive correlation: $V(p(t),x(t))\neq 0$ implies $p^{\top}(t)V(p(t),x(t))>0$, where $p(t)=v(t,x(t))$, for all $t\in[t_{0},T]$. $\square$
@@
$=\textstyle\sum_{i\in\mathbb{S}}\sum_{j\in\mathbb{S}}x_{j}(t)\dfrac{[v_{i}(t,x(t))-v_{j}(t,x(t))]_{+}^{2}}{q_{ji}(x(t))}.$
@@
Consider the modulo-addition stationary solution of the associated MFG system, as explained in (10). The stationary solution for (11) and (12) is given by: $x(t)=x^{\ast}$ and $v(t,x^{\ast})=\kappa(t-T)\mathbf{1}_{n}+F(x^{\ast}),~\forall t\in[t_{0},T]$, where $x^{\ast}$ is a Nash equilibrium of $F$. $\square$
@@
$\dot{v}(t,x^{\ast})=\kappa\mathbf{1}_{n}\implies v(t,x^{\ast})=\kappa(t-T)\mathbf{1}_{n}+v(T,x^{\ast}),$
with $v(T,x^{\ast})=F(x^{\ast})$. From (10b), it follows that $V(p(t),x^{\ast})=0$. By Definition 3, this condition implies $x_{i}^{\ast}[F_{j}(x^{\ast})-F_{i}(x^{\ast})]_{+}=0,~\forall i,j\in\mathbb{S}$, thereby establishing that $x^{\ast}$ is a Nash equilibrium.   $\blacksquare$
In the following theorem, we provide a convergence result for a certain class of contractive population games. The proof is provided in Appendix -B
Theorem 2
Suppose $q_{ij}(x),\,\forall i,j\in\mathbb{S}$ is continuously differentiable on $\Delta$, and a population game $F$ is contractive and admits an interior Nash equilibrium. Further, assume that $F$ satisfies
$\left(F(x)-F(y)\right)^{\top}(x-y)\leq-\epsilon\|x-y\|^{2},$ (13)
for a positive constant $\epsilon$. If $x(T/2)$ denotes the population state at time $t=T/2$, then as $T\to\infty$, we have $\|x(T/2)-x^{\ast}\|\to 0$, where $x^{\ast}$ is the Nash equilibrium of $F$. $\square$
@@
For any contractive population game $F$, modifying $F$ by defining $F_{\epsilon}(x)=F(x)-\epsilon\,(\ln(x_{1}+\delta)\,\cdots\,\ln(x_{n}+\delta))^{\top}$ with $\epsilon>0$, where $\delta>0$ is a sufficiently small constant. This modification ensures that $F_{\epsilon}$ satisfies (13) and has an interior Nash equilibrium. Let $x^{\ast}(\epsilon)$ denote the Nash equilibrium of $F_{\epsilon}$, which also serves as the $\epsilon$-Nash equilibrium of $F$. Since $F$ is a continuous function of $x$, it follows that:
@@
where $x^{\ast}$ is the Nash equilibrium of $F$. $\square$
@@ IV Illustrative Examples
Data: $F$, $x_{0}\in\Delta$, $t_{0},T\in\mathbb{R}_{+}$ with $T>t_{0}$, $N\in\mathbb{Z}_{+}$, $\varepsilon_{f}\in\mathbb{R}_{+}$
@@
10 $x^{(k+1)}_{[t_{0},T]}\leftarrow(1-a)x^{(k)}_{[t_{0},T]}+ax^{\prime}_{[t_{0},T]}$
11 $e\leftarrow\int_{t_{0}}^{T}\|x^{(k)}(t^{\prime})-x^{(k+1)}(t^{\prime})\|_{2}^{2}\,\mathrm{d}t^{\prime}$ $\triangleright$ To check if the fixed-point is reached
@@ 12 $k\leftarrow k+1$
We provide numerical examples to illustrate the main results of this paper. To compute the solutions for (11) and (12), we use Algorithm 1, adapted from the literature [Lauriere2023]. The algorithm is specifically designed to find the fixed point of (11) and (12).
@@
We note that Algorithm 1 requires solving two ordinary differential equations (Lines 4–5), whose computational cost depends on the length of the time horizon $T$ and the number of strategies $n$. In addition, decreasing the moving average weight $a$ (Line 6) can slow down convergence, thereby increasing the computational time, although it can also improve the accuracy and stability of the iterative process in the algorithm.
We consider the following two population games:
@@
$F(x)=\begin{pmatrix}F_{1}(x)\\ F_{2}(x)\\ F_{3}(x)\\ F_{4}(x)\\ F_{5}(x)\\ F_{6}(x)\end{pmatrix}=-\begin{pmatrix}2.5x_{1}+x_{2}\\ x_{1}+2.5x_{2}+x_{3}+0.5x_{5}\\ x_{2}+2.5x_{3}\\ 2.5x_{4}+x_{5}\\ 0.5x_{2}+x_{4}+2.5x_{5}+x_{6}\\ x_{5}+2.5x_{6}\end{pmatrix}$. (14)
For (4), we set $q_{ij}(x(t))=1,\forall i,j\in\mathbb{S}$, $t_{0}=0$, and $T=6$. We apply Algorithm 1 with $a=0.01$, $N=100$, and $\varepsilon_{f}=0$.
@@
For (4), we set $q_{ij}(x(t))=1,\forall i,j\in\mathbb{S}$, $t_{0}=0$, and $T=6$. We apply Algorithm 1 with $a=0.001$, $N=6000$, and $\varepsilon_{f}=0$.
@@
Fig. 3 compares the population state trajectories generated by the optimal strategy revision protocol with those obtained using the Smith ED model in the RPS game (15). As shown, the population state trajectory determined by the optimal protocol exhibits reduced oscillations and converges more quickly to the Nash equilibrium of (15).
Note that the congestion game (14) is a contractive game satisfying the assumptions of Theorem 2—namely, condition (13) and the existence of an interior Nash equilibrium. On the other hand, the RPS game (15) as a contractive game under Definition 2 but does not satisfy condition (13). Nevertheless, as illustrated in Fig. 3, our framework is still capable of designing the optimal revision protocol in this case, even without employing the technique discussed in Remark 2. This demonstrates its broader applicability to optimal design problems in PG.